Find the corresponding area between 𝑧 = 0 and each of the following.

1. z = 0.85

2. z = 1.27

3. z = 2.86

4. z = βˆ’1.05

5. z = βˆ’2.96


1
Expert's answer
2021-03-31T07:03:07-0400


S = |0.5 - f(z)|

According to z-score table:

1. S = |0.5 - 0.80324| = 0.30234

2. S = |0.5 - 0.89796| = 0.39796

3. S = |0.5 - 0.99788| = 0.49788

4. S = |0.5 - 0.29166| = 0.20834

5. S = |0.5 - 0.00154| = 0.49846